OpenStudy (anonymous):

Find the Mistake: provided below is a solution to the equation 4m − 8 = 2m. Choose the statement that is true about the given solution. Line # 4m − 8 = 2m 1 −2m − 8 = 0 2 −2m = 8 3 m = −4

OpenStudy (anonymous):

2) is right

OpenStudy (anonymous):

oh i didnt understand it

OpenStudy (anonymous):

4m-8=2m -8=2m-4m thus,-2m=-8

OpenStudy (anonymous):

ok so the whole problem was right

OpenStudy (anonymous):

yes and give me my medal

OpenStudy (anonymous):

already

OpenStudy (anonymous):

thanks and ask more